find the area bounded by the curve y=3x-x^3 and the line y=2


1
Expert's answer
2021-10-20T14:15:54-0400



Area enclosed:

"(3\times1)-\int_{-2}^{-\sqrt{3}}(3x-x^3)dx-\int^1_0(3x-x^3)dx+|\int^0_{-\sqrt{3}}(3x-x^3)dx|\\
\Rightarrow3-[\frac{3x^2}{2}-\frac{x^4}{4}]^{-\sqrt{3}}_{-2}-[\frac{3x^2}{2}-\frac{x^4}{4}]^{1}_{0}+|[\frac{3x^2}{2}-\frac{x^4}{4}]_{-\sqrt{3}}^{0}|\\
\Rightarrow3-[\frac{9}{2}-\frac{9}{4}-6+4]-[\frac{3}{2}-\frac{1}{4}-0+0]+|(0-0)-(\frac{9}{2}-\frac{9}{4})|\\
\Rightarrow 3-\frac{1}{4}-\frac{5}{4}+\frac{9}{4}\\
\Rightarrow \frac{15}{4} \ sq. units"